# Generalized Kummer surfaces over finite fields

### v3

@@ 5. Singularities of $A/G$
The cotangent space is a one-dimensional vector space over $k(\bar{\zeta}_{n})$. The Galois group of $k(\bar{\zeta}_{n})$ over $k$ is of order $2$ and the action on $\bar{\zeta}_{n}$ is given by the formula:
@@
Lemma 6.6.
Let $A$ be a non-ordinary surface, and let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to $p$. Assume that $\bar{\zeta}=\bar{\zeta}_{n}\in k$.
@@
    ()( ∗ ) $g(v_{1})=\zeta v_{1},\quad g(v_{2})=\zeta^{p}v_{2},\quad g(u_{1})=\zeta^{\prime}u_{1},\quad g(u_{2})=\zeta^{\prime p}u_{2},$
    where $\zeta^{\prime}$ is a prime to $n$ power of $\zeta$. If the action on $T_{0}^{*}(A)$ is symplectic, then there exists a basis such that ()(*)( ∗ ) holds with $\zeta^{\prime}=\zeta^{-1}$.
@@
    If $n\in\{3,4,6\}$, or $n\in\{8,12\}$, and $p\not\equiv\pm 1\bmod n$, then we can choose a basis of $M(A)\otimes{\mathbb{Q}}$ such that ()(*)( ∗ ) holds with $\zeta^{\prime}=\zeta^{-1}$, and
@@
Assume that $p^{2}\equiv 1\mod n$. Then the equality $\zeta^{p}=\zeta^{\prime}$ is equivalent to $(\zeta^{\prime})^{p}=\zeta$, because $(\zeta^{\prime})^{p}=\zeta^{p^{2}}=\zeta$. It follows that, if $\zeta^{p}=\zeta^{\prime}$, then $v_{1},u_{1}$ generate a Diedonné submodule of $M(A)$ of slope $1$. By duality, slopes of $A$ are 00 and $1$, i.e., $A$ is ordinary. A contradiction. Therefore, $\zeta^{p}\neq\zeta^{\prime}$, and $v_{1},v_{2},u_{1},u_{2}$ is a basis of $M(A)\otimes{\mathbb{Q}}$. Clearly, the action of $g$ on $T_{0}^{*}(A)$ is symplectic if and only if $\zeta^{\prime}=\zeta^{-1}$. Part $(1)$ is proved.
@@
Now, we construct a basis of $M(A)\otimes{\mathbb{Q}}$ that satisfy conditions of $(3)$; we call such a basis symplectic. Our assumptions imply that $p^{2}\equiv 1\mod n$. We already proved that in this case a basis $v_{1},v_{2},u_{1},u_{2}$ with property ()(*)( ∗ ) exists.
@@
Let $G=Q_{4n}$, and let $g\in G$ be a element of order $2n$. Assume that $2n$ is prime to $p$. Then over $k(\bar{\zeta}_{2n})$ there is a basis of $T_{0}^{*}(A)$ such that generators of $G$ act as matrices
@@
There exists an element $j\in G$ of order $4$ such that $jgj^{-1}=g^{-1}$. Therefore, $\bar{\zeta}_{2n}^{-1}$ is also an eigenvalue of $g$, and there exists a basis $v,u\in T^{*}(A)\otimes k(\bar{\zeta}_{2n})$ such that the matrix of $g$ is diagonal. It is straightforward to show that $j(v)=\gamma u$ for some $\gamma\in k(\bar{\zeta}_{2n})$, and $j(u)=-\gamma^{-1}v$. Therefore, the basis $v,\gamma u$ is a desired basis. ∎
@@
These matrices are symplectic; therefore, a symplectic basis for $Q_{8}$ is a symplectic basis for $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$, or $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ as well. ∎
@@
If there exists a supersingular abelian surface with a rigid action of a group $G$, then, according to Corollary 3.10, there exists a supersingular abelian surface $B$ over ${\mathbb{F}}_{p^{2}}$ with a rigid action of $G$. It follows that $G$ and $p$ satisfy the conditions in column $I$ by Theorem 4.3, and Proposition 6.1. Let $A$ be an abelian surface over $k$ with a rigid and symplectic action of $G$. Then the conditions in column $II$ follow from Lemmata 6.6 and 6.5.
Let us prove the existence part. According to Theorem 2.7, there exists a supersingular elliptic curve $E$ over ${\mathbb{F}}_{p^{2}}$ such that $\operatorname{End}^{\circ}(E)\cong{\mathbb{H}}_{p}$; it follows that $\operatorname{End}^{\circ}(E^{2})\cong M(2,{\mathbb{H}}_{p}).$ By Corollary 3.9, there exists a supersingular abelian surface with an action of a group $G$ if and only if there exists a homomorphism ${\mathbb{Q}}[G]^{\mathrm{rig}}\to M(2,{\mathbb{H}}_{p})$. According to Theorem 4.3, and Proposition 6.1, the conditions in the first column of the table imply that such a homomorphism exists.
Let $A=E^{2}$, and let ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$. We are going to prove that under the conditions of column $II$ there exists a submodule $M$ of $M(A)\otimes{\mathbb{Q}}$ such that the action on $M/FM$ is symplectic and $M\otimes{\mathbb{Q}}\cong M(A)\otimes{\mathbb{Q}}$; we will call such a submodule symplectic. Assume that a symplectic submodule exists. Then, according to Lemma 2.9, there exist an abelian variety $B$ and a $p$-isogeny $\varphi:A\to B$ such that $M(B)\cong M$; in particular, for all $\ell\neq p$ we have $T_{\ell}(A)\cong T_{\ell}(B)$. Therefore, the natural homomorphism $G\to\operatorname{End}^{\circ}(B)$ factors through $\operatorname{End}(B)$, and $G$ acts on $B$. By construction, the isogeny $\varphi$ is $G$–equivariant, and the action of $G$ on $B$ is rigid and symplectic.
@@
Assume that ${\mathbb{H}}={\mathbb{Q}}(\zeta_{n})$ is a quartic field extension of ${\mathbb{Q}}$, i.e., $n\in\{5,8,10,12\}$. Then either $n$ divides $p^{2}-1$, or ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$. In the first case, Lemma 6.6 implies that there exists a symplectic submodule $M$ in $M(A)$. If ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$, then according to Lemma 6.5 the action is symplectic if and only if $p^{2}\equiv-1\bmod 5$, i.e, $p\not\equiv\pm 1\bmod 5$.
@@
Let ${\mathbb{H}}$ be a central simple algebra over ${\mathbb{Q}}$, and let $A$ be an ordinary abelian variety over $k$ with a rigid action of a finite group $G$ such that the action factors through ${\mathbb{H}}:$
@@
Let $V$ be the irreducible representation of ${\mathbb{H}}$ over ${\mathbb{Q}}$. The induced representation of $G$ on $V$ is without fixed points, hence the characteristic polynomial of any element $g\in G$ of order $n$ is equal to some power of the cyclotomic polynomial $\Phi_{n}^{r_{g}}$. According to Proposition 2.12, $M(A)\cong M\oplus M^{*}$ is a direct sum of Diedonné modules. The action of $G$ on $M$ comes from the composition
@@
It follows that $M\otimes{\mathbb{Q}}_{p}$ is isomorphic to $(V\otimes{\mathbb{Q}}_{p})^{r_{M}}$ as a ptensor-${\mathbb{H}}\otimes{\mathbb{Q}}_{p}$-module, and the characteristic polynomial of the action of $g$ on $M\otimes{\mathbb{Q}}_{p}$ is equal to $\Phi_{n}^{r_{g}r_{M}}$. According to Proposition 2.12, we have the isomorphism $T^{*}(A)\cong M/pM$. The proposition is proved. ∎
@@
According to Corollary 3.9, there exists an abelian variety $B_{V}$ of dimension $d\dim B$ with a rigid action of $G$.
@@
    $G$ is a cyclic group of order $2,3,4$, or $6$;
@@
If $G$ is not cyclic, we will construct an ordinary elliptic curve $E$ such that there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]$ to $\operatorname{End}^{\circ}(E^{2})\cong M(2,L)$, where $L=\operatorname{End}^{\circ}(E)$. Then, according to Corollary 3.9, there exists an abelian surface $A$ isogenous to $E^{2}$ with a rigid action of $G$. If $G$ is $Q_{8},Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, then we use Proposition 6.9 with ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$ and find that the action on $A$ is symplectic.
Let $G$ be $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$. According to Corollary 2.4, if $L$ is not split in $2$, then there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]\cong{\mathbb{H}}_{2}$ to $M(2,L)$. We are going to apply Theorem 2.7 and find an elliptic curve $E$ such that its endomorphism algebra $L$ is not split in $2$. If $p\equiv 1\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 5\bmod 8$. If $p\equiv 3\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-4t+p$; the discriminant of this polynomial is $16-4p\equiv 5\bmod 8$. In both cases $L$ is inert in $2$, if $p>2$.
If $G=Q_{12}$, then ${\mathbb{Q}}^{\mathrm{rig}}[G]={\mathbb{H}}_{3}$. As before, we need an ordinary elliptic curve over ${\mathbb{F}}_{p}$ such that its endomorphism algebra $L$ is not split in $3$. If $p\equiv 1\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-3t+p$; the discriminant of this polynomial is $9-4p\equiv 2\bmod 3$. If $p\equiv 2\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 2\bmod 3$. In both cases $L$ is inert in $3$, if $p>3$. ∎
@@
The existence part follows from Theorems 6.3 and 6.11. On the other hand, according to Theorem 1.3, the action of $G$ on $A$ is rigid and symplectic. Since $p$ does not divide the order of $G$, it follows from Theorem 4.3 that the group $G$ belongs to the Katsura list from Theorem 1.4.
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then, according to Corollary 5.4, the abelian surface $A$ is supersingular, and ${\mathbb{F}}_{p^{2}}\subset k$. In this situation, according to Theorem 6.3, we have $p\not\equiv\pm 1\bmod n$. ∎
@@ 7.1. Zeta functions of abelian varieties and generalized Kummer surfaces
The zeta function of an abelian variety $A$ with the Weil polynomial $f_{A}=\prod_{j}(t-\pi_{j})$ can be computed as follows:
@@
$P_{i}(A,t)=\prod_{j_{1}<\dots<j_{i}}(1-t\pi_{j_{1}}\dots\pi_{j_{i}}).$
@@
$P_{2}(t)=\det(1-tF|\operatorname{NS}({\overline{X}})\otimes{\mathbb{Q}})=\prod_{r}\Phi_{r}(t)^{\lambda_{r}}$
@@
It follows from the definitions that the number of points on a supersingular $K3$ surface $X$ over ${\mathbb{F}}_{q}$ is equal to
@@
Let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with the Weil polynomial $f(t)=(t^{2}+\varepsilon q)^{2}$ with a rigid action of a group $G$, where $\varepsilon=\pm 1$. Denote the characteristic polynomial of the Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{G}$ by $h(t)$.
  • If $G$ is cyclic of order $3,4$, or $6$, then $h(t)=(t-q)^{2}(t+q)^{2}$.
@@
Let $V=H^{1}(\bar{A},{\mathbb{Q}}_{2})\otimes{\mathbb{Q}}(\sqrt{-\varepsilon q})$. Then $V=V_{1}\oplus V_{2}$, where $F$ acts as $(-1)^{s}\sqrt{-\varepsilon q}$ on $V_{s}$, and $s\in\{1,2\}$. Since the action of $G$ commute with the Frobenius action, both $V_{1}$ and $V_{2}$ are $G$-invariant, and
@@
Let $k={\mathbb{F}}_{q}$, where $q\equiv 1\bmod 4$. Then there exist a supersingular generalized Kummer surfaces $K(A,C_{4})$ over $k$ with the zeta functions $(1^{15},2^{7})$, and $(1^{20},2^{2})$. In both cases, the Weil polynomial of $A$ is equal to $(t^{2}-q)^{2}$.
@@
According to Corollary 3.9 and Lemma 6.5, there exists an isogeny from $A^{\prime}$ to an abelian surface $A$ with a rigid and symplectic action of $C_{4}$. Denote by $g\in\operatorname{End}(A)$ the image of a generator of $C_{4}$.
Since $F$ and $g$ commute, the vector space $V_{2}(A)=V_{1}\oplus V_{2}$ is a sum of $2$-dimensional $F$-invariant subspaces such that $g(V_{1})=V_{2}$. According to [Ry14, Proposition 3.5], there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T^{\prime}\subset V_{1}$ such that the action of $F$ on $T^{\prime}/2T^{\prime}$ is non-trivial, but the action of $F^{2}$ is trivial. Moreover, if $q\equiv 1\bmod 4$, then there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T\subset V_{1}$ such that $F$ acts on $T/2T\cong({\mathbb{Z}}/2{\mathbb{Z}})^{2}$ as identity.
By Lemma 2.5, there exists an abelian surface $B$ with a $2$-isogeny $A\to B$ and such that $T\oplus gT\cong T_{2}(B)$. In particular, there exists a rigid and symplectic action of $C_{4}$ on $B$. We have $B[2](k)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{16}$ and, according to Proposition 5.1, the singularities of $A/G$ are $4A_{3}+6A_{1}$; therefore, according to Lemma 7.2 and Proposition 5.5, the zeta function of the generalized Kummer surface $K(B,C_{4})$ is $(1^{20},2^{2})$.
In the same way, there exists an abelian surface $B^{\prime}$ with a rigid and symplectic action of $G$ such that $T^{\prime}\oplus gT^{\prime}\cong T_{2}(B^{\prime})$. A straightforward computation shows that on $B^{\prime}/G$ there are exactly two $F$-invariant singular points of type $A_{3}$ and two $F$-invariant singular points of type $A_{1}$. Moreover, since the action of $F^{2}$ is trivial on $B^{\prime}[2]({\bar{k}})$, other singular points are of degree $2$ over $k$. According to Proposition 5.5, there are $8$ $F$-invariant exceptional lines, and $5$ pairs of lines such that Frobenius acts non-trivially. It follows that, the zeta function of $K(B^{\prime},C_{4})$ is $(1^{15},2^{7})$. ∎
@@
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r}$, and $p>2$. There exists a generalized Kummer surface $X$ over $k$ corresponding to an abelian surface $A$ with the following parameters.

### v4

@@ 5. Singularities of $A/G$
The cotangent space is a one-dimensional vector space over $k(\bar{\zeta}_{n})$. The Galois group of $k(\bar{\zeta}_{n})$ over $k$ is of order 22 and the action on $\bar{\zeta}_{n}$ is given by the formula:
@@
Lemma $6.$
Let AA be a non-ordinary surface, and let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to pp. Assume that $\bar{\zeta}=\bar{\zeta}_{n}\in k$.
@@
    () $g(v_{1})=\zeta v_{1},\quad g(v_{2})=\zeta^{p}v_{2},\quad g(u_{1})=\zeta^{\prime}u_{1},\quad g(u_{2})=\zeta^{\prime p}u_{2},$
    where $\zeta^{\prime}$ is a prime to $n$ power of $\zeta$. If the action on $T_{0}^{*}(A)$ is symplectic, then there exists a basis such that ()(*) holds with $\zeta^{\prime}=\zeta^{-1}$.
@@
    If $n\in\{3,4,6\}$, or $n\in\{8,12\}$, and $p\not\equiv\pm 1\bmod n$, then we can choose a basis of $M(A)\otimes{\mathbb{Q}}$ such that ()(*) holds with $\zeta^{\prime}=\zeta^{-1}$, and
@@
Assume that $p^{2}\equiv 1\mod n$. Then the equality $\zeta^{p}=\zeta^{\prime}$ is equivalent to $(\zeta^{\prime})^{p}=\zeta$, because $(\zeta^{\prime})^{p}=\zeta^{p^{2}}=\zeta$. It follows that, if $\zeta^{p}=\zeta^{\prime}$, then $v_{1},u_{1}$ generate a Diedonné submodule of $M(A)$ of slope 11. By duality, slopes of AA are 0 and 11, i.e., AA is ordinary. A contradiction. Therefore, $\zeta^{p}\neq\zeta^{\prime}$, and $v_{1},v_{2},u_{1},u_{2}$ is a basis of $M(A)\otimes{\mathbb{Q}}$. Clearly, the action of $g$ on $T_{0}^{*}(A)$ is symplectic if and only if $\zeta^{\prime}=\zeta^{-1}$. Part $(1)$ is proved.
@@
Now, we construct a basis of $M(A)\otimes{\mathbb{Q}}$ that satisfy conditions of $(3)$; we call such a basis symplectic. Our assumptions imply that $p^{2}\equiv 1\mod n$. We already proved that in this case a basis $v_{1},v_{2},u_{1},u_{2}$ with property ()(*) exists.
@@
Let $G=Q_{4n}$, and let $g\in G$ be a element of order $2n$. Assume that $2n$ is prime to pp. Then over $k(\bar{\zeta}_{2n})$ there is a basis of $T_{0}^{*}(A)$ such that generators of $G$ act as matrices
@@
There exists an element $j\in G$ of order 44 such that $jgj^{-1}=g^{-1}$. Therefore, $\bar{\zeta}_{2n}^{-1}$ is also an eigenvalue of $g$, and there exists a basis $v,u\in T^{*}(A)\otimes k(\bar{\zeta}_{2n})$ such that the matrix of $g$ is diagonal. It is straightforward to show that $j(v)=\gamma u$ for some $\gamma\in k(\bar{\zeta}_{2n})$, and $j(u)=-\gamma^{-1}v$. Therefore, the basis $v,\gamma u$ is a desired basis. ∎
@@
These matrices are symplectic; therefore, a symplectic basis for $Q_{8}$ is also a symplectic basis for $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ and $\operatorname{SL}_{2}({\mathbb{F}}_{5})$. ∎
@@
If there exists a supersingular abelian surface with a rigid action of a group $G$, then, according to Corollary 3.9, there exists a supersingular abelian surface $B$ over ${\mathbb{F}}_{p^{2}}$ with a rigid action of $G$. It follows that $G$ and $p$ satisfy the conditions in column II by Theorem 4.3, and Proposition 6.1. Let AA be an abelian surface over $k$ with a rigid and symplectic action of $G$. Then the conditions in column IIII follow from Lemmata 6.6 and 6.5.
Let us prove the existence part. According to Theorem 2.7, there exists a supersingular elliptic curve $E$ over ${\mathbb{F}}_{p^{2}}$ such that $\operatorname{End}^{\circ}(E)\cong{\mathbb{H}}_{p}$; it follows that $\operatorname{End}^{\circ}(E^{2})\cong M(2,{\mathbb{H}}_{p}).$ By Theorem 1.3, there exists a supersingular abelian surface with an action of a group $G$ if and only if there exists a homomorphism ${\mathbb{Q}}[G]^{\mathrm{rig}}\to M(2,{\mathbb{H}}_{p})$. According to Theorem 4.3, and Proposition 6.1, the conditions in the first column of the table imply that such a homomorphism exists.
Let $A=E^{2}$, and let ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$. We are going to prove that under the conditions of column IIII there exists a submodule $M$ of $M(A)\otimes{\mathbb{Q}}$ such that the action on $M/FM$ is symplectic and $M\otimes{\mathbb{Q}}\cong M(A)\otimes{\mathbb{Q}}$; we will call such a submodule symplectic. Assume that a symplectic submodule exists. Then, according to Lemma 2.9, there exist an abelian variety $B$ and a $p$-isogeny $\varphi:A\to B$ such that $M(B)\cong M$; in particular, for all $\ell\neq p$ we have $T_{\ell}(A)\cong T_{\ell}(B)$. Therefore, the natural homomorphism $G\to\operatorname{End}^{\circ}(B)$ factors through $\operatorname{End}(B)$, and $G$ acts on $B$. By construction, the isogeny $\varphi$ is GG–equivariant, and the action of $G$ on $B$ is rigid and symplectic.
@@
Assume that ${\mathbb{H}}={\mathbb{Q}}(\zeta_{n})$ is a quartic field extension of $\mathbb{Q}$}, that is, $n\in\{5,8,10,12\}$. Then either $n$ divides $p^{2}-1$, or ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$. In the first case, Lemma 6.6 implies that there exists a symplectic submodule $M$ in $M(A)$. If ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$, then according to Lemma 6.5 the action is symplectic if and only if $p^{2}\equiv-1\bmod 5$, that is, $p\not\equiv\pm 1\bmod 5$.
@@
Let $\mathbb{H}$} be a central simple algebra over $\mathbb{Q}$}, and let AA be an ordinary abelian variety over $k$ with a rigid action of a finite group $G$ such that the action factors through $\mathbb{H}$}:
@@
Let $V$ be the irreducible representation of $\mathbb{H}$} over $\mathbb{Q}$}. The induced representation of $G$ on $V$ is without fixed points; therefore, the characteristic polynomial of any element $g\in G$ of order $n$ is equal to some power of the cyclotomic polynomial $\Phi_{n}^{r_{g}}$. According to Proposition 2.12, $M(A)\cong M\oplus M^{*}$ is a direct sum of Diedonné modules. The action of $G$ on $M$ comes from the composition
@@
It follows that $M\otimes{\mathbb{Q}}_{p}$ is isomorphic to $(V\otimes{\mathbb{Q}}_{p})^{r_{M}}$ as a ${\mathbb{H}}\otimes{\mathbb{Q}}_{p}$-module, and the characteristic polynomial of the action of $g$ on $M\otimes{\mathbb{Q}}_{p}$ is equal to $\Phi_{n}^{r_{g}r_{M}}$. According to Proposition 2.12, we have the isomorphism $T^{*}(A)\cong M/pM$. The proposition is proved. ∎
@@
According to Theorem 1.3, there exists an abelian variety $B_{V}$ of dimension $d\dim B$ with a rigid action of $G$.
@@
    $G$ is a cyclic group of order $2,3,4$, or 66;
@@
If $G$ is not cyclic, we will construct an ordinary elliptic curve $E$ such that there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]$ to $\operatorname{End}^{\circ}(E^{2})\cong M(2,L)$, where $L=\operatorname{End}^{\circ}(E)$. Then, according to Theorem 1.3, in the isogeny class of $E^{2}$ there exists an abelian surface AA with a rigid action of $G$. If $G$ is $Q_{8},Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, then we use Proposition 6.9 with ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$ and find that the action on AA is symplectic.
Let $G$ be $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$. According to Corollary 2.4, if $L$ is not split at 22, then there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]\cong{\mathbb{H}}_{2}$ to $M(2,L)$. We are going to apply Theorem 2.7 and find an elliptic curve $E$ such that its endomorphism algebra $L$ is not split at 22. If $p\equiv 1\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 5\bmod 8$. If $p\equiv 3\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-4t+p$; the discriminant of this polynomial is $16-4p\equiv 5\bmod 8$. In both cases $L$ is inert at 22, if $p>2$.
If $G=Q_{12}$, then ${\mathbb{Q}}^{\mathrm{rig}}[G]={\mathbb{H}}_{3}$. As before, we need an ordinary elliptic curve over ${\mathbb{F}}_{p}$ such that its endomorphism algebra $L$ is not split at 33. If $p\equiv 1\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-3t+p$; the discriminant of this polynomial is $9-4p\equiv 2\bmod 3$. If $p\equiv 2\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 2\bmod 3$. In both cases $L$ is inert at 33, if $p>3$. ∎
@@
The existence part follows from Theorems 6.3 and 6.11. On the other hand, according to Theorem 1.4, the action of $G$ on AA is rigid and symplectic. Since $p$ does not divide the order of $G$, it follows from Theorem 4.3 that the group $G$ belongs to the Katsura list from Theorem 1.5.
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then, according to Corollary 5.4, the abelian surface AA is supersingular, and ${\mathbb{F}}_{p^{2}}\subset k$. In this situation, according to Theorem 6.3, we have $p\not\equiv\pm 1\bmod n$. ∎
@@ 7.1. Zeta functions of abelian varieties and generalized Kummer surfaces
The zeta function of an abelian variety AA with the Weil polynomial $f_{A}=\prod_{j}(t-\pi_{j})$ can be computed as follows:
@@
$P_{i}(A,t)=\prod_{j_{1}<\dots<j_{i}}(1-\pi_{j_{1}}\dots\pi_{j_{i}}t).$
@@
$P_{2}(t)=\det(1-Ft|\operatorname{NS}({\overline{X}})\otimes{\mathbb{Q}})=\prod_{r}\Phi_{r}(t)^{\lambda_{r}}$
@@
From the definitions it follows that the number of points on a supersingular $K3$ surface $X$ over ${\mathbb{F}}_{q}$ is equal to
@@
Let AA be an abelian surface over ${\mathbb{F}}_{q}$ with the Weil polynomial $f(t)=(t^{2}+\varepsilon q)^{2}$ with a rigid action of a group $G$, where $\varepsilon=\pm 1$. Denote the characteristic polynomial of the Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{G}$ by $h(t)$.
  • If $G$ is cyclic of order $3,4$, or 66, then $h(t)=(t-q)^{2}(t+q)^{2}$.
@@
Let $V=H^{1}(\bar{A},{\mathbb{Q}}_{2})\otimes{\mathbb{Q}}(\sqrt{-\varepsilon q})$. Then $V=V_{1}\oplus V_{2}$, where $F$ acts as $(-1)^{s}\sqrt{-\varepsilon q}$ on $V_{s}$, and $s\in\{1,2\}$. Since the action of $G$ commutes with the Frobenius action, both $V_{1}$ and $V_{2}$ are $G$-invariant, and
@@
Assume that $q\equiv 1\bmod 4$. Then there exist supersingular generalized Kummer surfaces $K(A,C_{4})$ over ${\mathbb{F}}_{q}$ with zeta functions $(1^{15},2^{7})$ and $(1^{20},2^{2})$. In both cases, the Weil polynomial of AA is equal to $(t^{2}-q)^{2}$.
@@
According to Theorem 1.3 and Lemma 6.5, there exists an isogeny from $A^{\prime}$ to an abelian surface AA with a rigid and symplectic action of $C_{4}$. Denote by $g\in\operatorname{End}(A)$ the image of a generator of $C_{4}$.
Since $F$ and $g$ commute, the vector space $V_{2}(A)=V_{1}\oplus V_{2}$ is a sum of 22-dimensional $F$-invariant subspaces such that $g(V_{1})=V_{2}$. According to [Ry14, Proposition 3.5], there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T^{\prime}\subset V_{1}$ such that the action of $F$ on $T^{\prime}/2T^{\prime}$ is non-trivial, but the action of $F^{2}$ is trivial. Moreover, if $q\equiv 1\bmod 4$, then there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T\subset V_{1}$ such that $F$ acts on $T/2T\cong({\mathbb{Z}}/2{\mathbb{Z}})^{2}$ as identity.
By Lemma 2.5, there exists an abelian surface $B$ with a 22-isogeny $A\to B$ and such that $T\oplus gT\cong T_{2}(B)$. In particular, there exists a rigid and symplectic action of $C_{4}$ on $B$. We have $B[2]({\mathbb{F}}_{q})\cong({\mathbb{Z}}/2{\mathbb{Z}})^{16}$ and, according to Proposition 5.1, the singularities of $A/G$ are $4A_{3}+6A_{1}$; therefore, since $q\equiv 1\bmod 4$, according to Lemma 7.2 and Proposition 5.5, the zeta function of the generalized Kummer surface $K(B,C_{4})$ is $(1^{20},2^{2})$.
In the same way, there exists an abelian surface $B^{\prime}$ with a rigid and symplectic action of $G$ such that $T^{\prime}\oplus gT^{\prime}\cong T_{2}(B^{\prime})$. A straightforward computation shows that on $B^{\prime}/G$ there are exactly two $F$-invariant singular points of type $A_{3}$ and two $F$-invariant singular points of type $A_{1}$. Moreover, since the action of $F^{2}$ is trivial on $B^{\prime}[2]$, other singular points are of degree 22 over ${\mathbb{F}}_{q}$. According to Proposition 5.5, there are 88 $F$-invariant exceptional lines, and 55 pairs of lines such that Frobenius acts non-trivially. It follows that the zeta function of $K(B^{\prime},C_{4})$ is $(1^{15},2^{7})$. ∎
@@
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r}$, and $p>2$. There exists a generalized Kummer surface $X$ over $k$ that corresponds to an abelian surface AA with the following parameters.